OpenStudy (anonymous):

Find the formula for a linear function given in a table x h(x) 3 0 4 6 5 12

OpenStudy (anonymous):

Using any 2 points, lets take (3,0) and (4,6) Find the gradient/slope (6-0)/ (4-3) = 6 y=6x+c where c is a constant Substituting (3,0), 0= 6(3) +c c= -18 Eqn of line is y= 6x -18

OpenStudy (anonymous):

ok thank you